p(x)=x4+5x3+4x2-3x+9=(x-k)2(ax2+bx+c) where k is an integer and a,b,c are constants
find the value of 5a-7b+8c

Grade:10

1 Answers

Sher Mohammad IIT Delhi
9 years ago
for x=0 . 9=-2kc, let c=3/2 imply k=-3
x=1 16=(1-k)2(a+b+c)
16=4*2(a+b+c)
a+b+c=2
a+b= .5
for x=-1
12= 2*2(a-b+c)
a-b=3-1.5=1.5
_ a=1
b=-.5